The problem is to foil [4(x+2)^.5][1+(x+2)^-.5]=? or how to simplify this problem?
Answer by Tatiana_Stebko(1539)

[4(x+2)^.5][1+(x+2)^-.5]=4(x+2)^.5*1+4(x+2)^.5(x+2)^-.5=4(x+2)^.5+4(x+2)^(.5-.5)=4(x+2)^.5+4(x+2)^0=4(x+2)^.5+4